OpenStudy (swag):

Write the following value in standard notation. 3.89 x 104 389,000 38,900,000 389,410 38,900

OpenStudy (swag):

@jazy

OpenStudy (anonymous):

10^4 = 10 * 10 * 10 * 10 = 10,000 3.89 * 10,000 = 38,900